# Learning What’s Missing: Attention Dispersion and EMA Stabilization in Length Generalization

### v1

@@ Abstract
We study length generalization in transformers through the set complement task, where a model must predict a uniform distribution over tokens absent from an input sequence—an ability central to board-game style reasoning. Our main theoretical result establishes two statements. First, we prove tight bounds on embedding and value dimensions for single-layer attention-only transformers. Second, we show that if such a model achieves balanced logit displacement at lengths 1 and 2, then it must generalize to longer sequences, though with reduced precision. A mechanistic reading of the proof explains this limitation: as more tokens are attended to, softmax compresses logit displacements, eroding separation between valid and invalid outputs. Training dynamics also suggest a second obstacle: when many next tokens are possible, updates become noisy. We hypothesize that dropout can counteract the first effect and Exponential Moving Average (EMA) the second. We validate these hypotheses through random hyperparameter search on the set complement task, which confirms both mechanisms. We then test OthelloGPT, a GPT-1 style model trained on random Othello moves, and find that EMA again improves length generalization in this more complex setting.
@@ 1 Introduction
Since it was discovered that transformer [1]-based large language models (LLMs) can be aligned with human interests [2], LLM agents are being deployed in roles as diverse as application developers, counsellors, job interviewers, or research assistants. For safety and efficiency both it is paramount that we understand how these agents make their decisions.
Prior work revealed that to improve the reasoning capabilities of transformers, one can prompt [3] or train them to think in small steps. This makes reasoning transformers produce their answers akin to game-playing agents that map out various potential trajectories before making a move.
@@
We abstract this task as the Set Complement Task, introduced in Subsection 3.1: given an input sequence of tokens without repetition, the model has to output a uniform distribution on the tokens absent from the input sequence. Note that we aim for more that top-1 accuracy, that is for the model to predict as most probable next token one that is missing from the input: such a basic component has to be learnt free from bias.
@@
This connects our work to the topic of length generalization: if a model robustly learned to perform an algorithmic task, then it should be able to produce a correct output on input sequences longer than those in its training set. It is an active field of study which tasks can transformers length generalize on: we discuss this in detail in Section 2.
It is of particular interest in the study of length generalization how to overcome obstacles to it. In Subsection 4.3, a mechanistic reading of the way our models make their inferences identifies the reduction in precision as a particular case of attention dispersion [4, 5]: softmax attention reduces displacements between attention weights as sequence length increases. We hypothesize that increased dropout may be able to increase the amplitude of value vectors and thus mitigate this effect.
@@
We investigate the effect of our proposed strategies in a random hyperparameter search experiment, the results of which are reported in Section 5. We find that both our strategies reliably improve performance metrics.
To validate our methods in a more complex setting, in Section 6, we study length generalization in OthelloGPT [6, 7]. In this setup, we train GPT-1 style models to predict legal moves in random Othello games. We find that BEMA robustly improves performance metrics in this case too.
To summarize, our contributions are as follows:
  1. 1.
    We introduce the Set Complement Task, a simple algorithmic task that any board game playing agent should be able to solve and which can serve as a benchmark for length generalization and training with noisy gradients.
  2. 2.
    We characterize the minimal transformers that can solve the task. In particular, we prove a length generalization property: if a model can solve the task for input sequences of length 1 and 2 in a balanced way, then it must length generalize, albeit with reduced precision.
  3. 3.
    We mechanistically analyze inference and training dynamics, pointing out attention dispersion and noisy gradients as potential obstacles for length generalization. We propose increased dropout and BEMA as respective mitigations.
  4. 4.
    We conduct a random hyperparameter search experiment and show that our proposed strategies reliably improve performance metrics.
  5. 5.
    We study how our methods generalize to length generalization in OthelloGPT. We show that BEMA robustly improves performance in this setting too.
@@ 2 Related Work
Mechanistic Interpretability aims to find minimal subnetworks, so-called circuits, of an artificial neural network that satisfy a given task. In the case of transformers, several such circuits have been identified such as induction heads [8], indirect object identification circuits [9], greater-than circuits [10], and retrieval heads [11]. Of particular interest to the present work are the studies on OthelloGPT, which showed that in the residual stream of a GPT-1 style model trained to predict legal moves on random Othello games, via nonlinear [6] and linear [7] probing, one can find representations of board state. We intend to extend the compendium of known circuits by minimal transformers that can solve the Set Completion Task.
@@
Length Generalization studies the conditions under which sequence-to-sequence models retain their performance on inputs longer than those seen during training. One train of results seeks to find criteria for algorithms transformers can length generalize on [12]. An important theoretical tool in this direction is the Restricted Access Sequence Processing Language (RASP) [13], a programming language that a transformer can implement. It was conjectured that length generalization is possible if there is a simple implementation in RASP-L [14]. Afterwards, a version of this conjecture was proven [15] using limit transformers, and a version of the C-RASP language [16]. In the usual algorithmic approach to the study of length generalization, if multiple solutions are possible, then the model has to predict the set of valid solutions as a singleton. We aim to bring in an alternative point of view closer to the spirit of language modeling: if multiple solutions are possible, the model should output a uniform distribution between them.
Attention Dispersion [4, 5] is a drawback of softmax attention that forbids generalization to arbitrary sequence lengths both in toy and language models. Most proposed solutions, such as adjusting attention logits [17] or next token logit temperatures [5] by a sequence length-dependent terms, make a change to the most commonly used architectures. In the present work, we are interested in whether we can use solutions that are still in line with mainstream models.
@@
The study of the next token distributions output by LLMs brings a detailed view on how the models generate theirs answers, and how expressive they can get. An important part of this point of view is how calibrated are the models, that is how well do the predicted next token distributions approximate the target next token distribution. It is shown [18] that pretrained models as small as GPT-Neo-1.3B surpass humans in next token prediction on the OpenWebText dataset [19], both in top-1 accuracy, and perplexity. However, calibration to the pretraining corpus can be proven to bring in hallucinations [20], at the very least on facts not present in the training dataset, given the assumption that there are exponentially more ways to complete a sentence in an untruthful way. Neither base or aligned models are calibrated in numeric contexts such as generating tokens from a uniform distribution [21], rather they have strong systematic biases such as dependence on token order. Through soft [22] and hard [23] prompt tuning experiments, it was discovered [24] that transformers are more capable of outputting distributions of very low or very high entropy, those with outliers, or those that were output by other transformers. The experiments were conducted both on pretrained and randomly initialized models, thus indicating that the limits in expressivity may stem from the transformer architecture, or the softmax output. In our work, we also investigate if the model learns the uniform distribution among valid next tokens, thus indicating that predictions are free from bias.
@@ 3.2 Minimal Transformers
The next token logit bias matrix $\mathbf{B}=\mathbf{E}\mathbf{U}\in\mathbb{R}^{v\times v}$ is the product of the token embedding parameter matrix $\mathbf{E}\in\mathbb{R}^{v\times d}$ and the unembedding parameter matrix $\mathbf{U}\in\mathbb{R}^{d\times v}$. For any $1\leq t\leq v$, the $t$-th row $\mathbf{E}_{t,:}\in\mathbb{R}^{d}$ is a $d$-dimensional trainable vector the token $t\in\mathbb{V}$ is mapped to. We call $d$ the embedding dimension, the vector space $\mathbb{R}^{d}$ the residual stream, and the vector space $\mathbb{R}^{v}$ the logit space. In our minimal transformers, we do not use positional encodings. During training, dropout [25] is optionally applied to the token embedding vectors. We call this dropout operation the embedding dropout.
The unnormalized per-token attention weight matrix $\mathbf{A}$ is formed as follows: First, RMSNorm [26] is applied to the token embedding vectors. We denote by $\mathbf{E}^{\prime}$ the token embedding parameter matrix with RMSNorm applied to its rows. Then, we form the query and key per-token matrices $\mathbf{Q}=\mathbf{E}^{\prime}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{E}^{\prime}\mathbf{W}_{K}$ via the query and key parameters matrices $\mathbf{W}_{Q},\mathbf{W}_{K}\in\mathbb{R}^{d\times d_{k}}$. We call $d_{k}$ the key dimension. In our setup, the key dimension is not necessarily equal to the embedding dimension divided by the number of attention heads. The per-token attention logit matrix is the product $\mathbf{A}^{\prime}=\mathbf{Q}\mathbf{K}^{T}$. This yields the unnormalized per-token attention weight matrix via the elementwise formula: $a_{i,j}=\exp(a^{\prime}_{i,j}/\sqrt{d_{k}})$. Note that as we define the output in Equation (3) for one input sequence only, we do not have to be explicit about causal attention. As we found it to be detrimental in initial experiments on the set complement task, we do not use attention dropout.
The next token logit displacement matrix $\mathbf{D}=\mathbf{E}^{\prime}\mathbf{W}_{V}\mathbf{W}_{O}\mathbf{U}$ is formed via the value and output parameter matrices $\mathbf{W}_{V}\in\mathbb{R}^{d\times d_{v}}$, $\mathbf{W}_{O}\in\mathbb{R}^{d_{v}\times d}$. We call $d_{v}$ the value dimension. In our setup, the value dimension is not necessarily equal to the key dimension, nor is it necessarily equal to the embedding dimension divided by the number of attention heads. During training, dropout is optionally applied to the rows of the matrix product $\mathbf{E}^{\prime}\mathbf{W}_{V}\mathbf{W}_{O}$ before multiplication from the right by the unembedding parameter matrix $\mathbf{U}$. We call this dropout operation the residual dropout.
@@ 4.1 A Hardcoded, Minimal Solution
For any vocabulary size $v$, we now give a formula for a model that is arbitrarily close to being perfect if we choose $C>0$ large enough. It uses embedding and value dimensions $d=d_{v}=v-1$, and key dimension $d_{k}=1$. We can use the following parameter matrices:
$\mathbf{E}=\begin{pmatrix}1&0&\dotsb&0\\ 0&1&\dotsb&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dotsb&1\\ -1&-1&\dotsb&-1\end{pmatrix},\,\mathbf{U}=\begin{pmatrix}-1&0&\dotsb&0&0\\ 0&-1&\dotsb&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dotsb&-1&0\end{pmatrix}$,
$\mathbf{W}_{Q}=\mathbf{W}_{K}=\boldsymbol{0},\,\mathbf{W}_{V}=vC\mathbf{I},\,\mathbf{W}_{O}=I.$
As the embedding vectors have constant length, we can set the RMSNorm scaling parameters to get $\mathbf{E}^{\prime}=\mathbf{E}$. Note that with these parameters, we get
@@
Note that the hardcoded model $f_{\theta}$ has constant attention. Since when training transformers with weight decay induces low-rank attention logit matrices [27], we will continue our theoretical investigation with the assumption of constant attention. Therefore, the formula (3) for the next token logit vector $f_{\theta}(\mathbf{t})$ at the input sequence $\mathbf{t}=(t_{1},\dotsc,t_{s})\in\mathbb{X}$ simplifies to the following:

### v2

# Post-Norm can Resharpen Attention

@@ Abstract
Length Generalization is the essential capacity of autonomous agents to perform tasks in longer contexts than those encountered during training. To systematically study this feat, we test how well models can approximate the next token distributions in algorithmic tasks. This is to take into account the realistic possibility of multiple next tokens being legal. We present a prototypical benchmark for this line of study: in the Set Complement Task, the model needs to output a uniform distribution over tokens not in the input. We prove a theorem that states simple transformers can length generalize on this task, however, with performance degradation due to attention dispersion. A mechanistic reading of how dispersion takes effect lets us discover a remedy: Post-Norm can Resharpen Attention. We present experimental evidence to support this idea. We also show that Exponential Moving Averages can help the issue of noisy gradients that arises when many next tokens are legal. We validate the general applicability of our proposed methods on a suite of formal language experiments. Our source code will be available upon publication.
attention dispersion, length generalization, post-norm, transformers
@@ 1 Introduction
Since it was discovered that transformer (Vaswani et al., 2017)-based large language models (LLMs) can be aligned with human interests (Ouyang et al., 2022), LLM agents are being deployed in roles as diverse as application developers, counsellors, job interviewers, or research assistants. For both safety and efficiency it is paramount that we understand how these agents make their decisions.
Prior work revealed that to improve the reasoning capabilities of transformers, one can prompt (Wei et al., 2022) or train them to think in small steps. This makes reasoning transformers produce their answers akin to game-playing agents that map out various potential trajectories before making a move.
@@
We abstract this task as the Set Complement Task, introduced in Subsection 3.1: given an input sequence of tokens without repetition, the model has to output a uniform distribution on the tokens absent from the input sequence. Note that we aim for more that top-1 accuracy, that is for the model to predict as most probable next token one that is missing from the input: such a basic component has to be learnt free from bias, that is, without predicting one valid token more often than the other.
@@
This connects our work to the topic of length generalization: if a model robustly learned to perform an algorithmic task, then it should be able to produce a correct output on input sequences longer than those in its training set. It is an active field of study which tasks transformers can length generalize on: we discuss this in detail in Section 2.
Overcoming obstacles to length generation is of particular interest for this line of study. In Subsection 4.3, a mechanistic reading of the way our models make their inferences identifies the reduction in precision as a particular case of attention dispersion (Hahn, 2020; Veličković et al., 2025): by taking a mean along the sequential dimention, softmax attention reduces the relative differences between attention weights as sequence length increases. We hypothesize that normalization of the output of the attention blocks may be able to undo the reduction in the amplitude of value vectors and thus mitigate this effect.
@@
We investigate the effect of our proposed strategies in random hyperparameter search experiments. We describe our experimental protocol in Section 5. We report our experiment results in Section 6: in Subsection 6.1, we provide experimental evidence that our proposed methods indeed foster length generalization in the Set Complement Task. We show that our methods have an effect in further length generalization tasks: Maximum Retrieval in Subsection 6.2, a bounded Dyck language in Subsection 6.3, and Tomita languages 3 and 7 in 6.4.
@@ 2 Related Work
Length Generalization studies the conditions under which sequence-to-sequence models retain their performance on inputs longer than those seen during training. One train of results seeks to find criteria for algorithms transformers can length generalize on (Bhattamishra et al., 2020). An important theoretical tool in this direction is the Restricted Access Sequence Processing Language (RASP) (Rush & Weiss, 2023), a programming language that a transformer can implement. It was conjectured that length generalization is possible if there is a simple implementation in RASP-L (Zhou et al., 2024). Afterwards, a version of this conjecture was proven (Huang et al., 2025) using limit transformers, and a version of the C-RASP language (Yang & Chiang, 2024). In the usual algorithmic approach to the study of length generalization, if multiple solutions are possible, then the model has to predict the set of valid solutions as a singleton. We aim to bring in an alternative point of view closer to the spirit of language modeling: if multiple solutions are possible, the model should learn to output the correct distribution between them.
Attention Dispersion (Hahn, 2020; Veličković et al., 2025) is a drawback of softmax attention that forbids generalization to arbitrary sequence lengths both in toy and language models. Multiple previous works considered fixing this dilution by an explicit, length-dependent transformation of the attention formula, either a $\log(s)$ multiplier on the attention logits (Chiang & Cholak, 2022), or a fixed temperature function found by polynomial interpolation (Veličković et al., 2025; Peng et al., 2024). However, to our knowledge, in the context of length generalization, post-normalization was only considered theoretically (Hahn, 2020) or in custom-made constructions (Yao et al., 2021), not as a general helper.
In the original transformer (Vaswani et al., 2017), normalization layers are put on the residual stream after the attention and the feedforward block updates have been applied. This approach is called called post-normalization. Later on, it was discovered that this makes training unstable at initialization (Xiong et al., 2020), so normalizing the inputs to the attention and feedforward blocks was proposed, called pre-normalization. However, it was observed that pre-normalization can cause massive activations to appear later in training (Sun et al., 2024). Further options were considered in specific architectures, for example: normalizing both the input and the output of the residual blocks as Sandwich-LN in the text-to-image model Cogview (Ding et al., 2021), and normalizing the output of the residual connections in the language models Swin Transformer 2 V2 (Liu et al., 2022). These ideas were combined to peri-normalization (Kim et al., 2025), where the inputs and outputs of the residual blocks, the input of the unembedding, and optionally the output of the embedding are normalized. In all these works, normalization is studied from the point of view of stable training, not length generalization.
Mechanistic Interpretability aims to find minimal subnetworks, so-called circuits, of an artificial neural network that satisfy a given task. In the case of transformers, several such circuits have been identified such as induction heads (Elhage et al., 2021), indirect object identification circuits (Wang et al., 2023), greater-than circuits (Hanna et al., 2023), and retrieval heads (Wu et al., 2025). Of particular interest to the present work are the studies on OthelloGPT, which showed that in the residual stream of a GPT-1 style model trained to predict legal moves on random Othello games, via nonlinear (Li et al., 2023) and linear (Nanda et al., 2023) probing, one can find representations of board state. We intend to extend the compendium of known circuits by minimal transformers that can solve the Set Completion Task.
@@
The study of the next token distributions output by LLMs brings a detailed view on how the models generate theirs answers, and how expressive they can get. An important part of this point of view is how calibrated are the models, that is how well do the predicted next token distributions approximate the target next token distribution. It is shown (Shlegeris et al., 2024) that pretrained models as small as GPT-Neo-1.3B surpass humans in next token prediction on the OpenWebText dataset (Gokaslan & Cohen, 2019), both in top-1 accuracy, and perplexity. However, calibration to the pretraining corpus can be proven to bring in hallucinations (Kalai & Vempala, 2024), at the very least on facts not present in the training dataset, given the assumption that there are exponentially more ways to complete a sentence in an untruthful way. Neither base or aligned models are calibrated in numeric contexts such as generating tokens from a uniform distribution (Lovering et al., 2025), rather they have strong systematic biases such as dependence on token order. Through soft (Li & Liang, 2021) and hard (Wallace et al., 2019) prompt tuning experiments, it was discovered (Wang et al., 2025) that transformers are more capable of outputting distributions of very low or very high entropy, those with outliers, or those that were output by other transformers. The experiments were conducted both on pretrained and randomly initialized models, thus indicating that the limits in expressivity may stem from the transformer architecture, or the softmax output. In our work, we also investigate if the model learns the true distribution among next tokens, thus indicating that predictions are free from bias.
@@ 3 Preliminaries
In the first, theoretical part of the paper, we will study single layer, attention only transformers trained on the Set Complement Task. In this Section, we shall introduce the requisite notions. In the second half of the paper, we will report on experiments on full transformers trained on more complex tasks.
@@ 3.2 Minimal Transformers
The next token logit bias matrix $\mathbf{B}=\mathbf{E}\mathbf{U}\in\mathbb{R}^{v\times v}$ is the product of the token embedding parameter matrix $\mathbf{E}\in\mathbb{R}^{v\times d}$ and the unembedding parameter matrix $\mathbf{U}\in\mathbb{R}^{d\times v}$. For any $1\leq t\leq v$, the $t$-th row $\mathbf{E}_{t,:}\in\mathbb{R}^{d}$ is a $d$-dimensional trainable vector the token $t\in\mathbb{V}$ is mapped to. We call $d$ the embedding dimension, the vector space $\mathbb{R}^{d}$ the residual stream, and the vector space $\mathbb{R}^{v}$ the logit space. In our minimal transformers, we do not use positional encodings.
The unnormalized per-token attention weight matrix $\mathbf{A}$ is formed as follows: We get the query and key per-token matrices $\mathbf{Q}=\mathbf{E}^{\prime}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{E}^{\prime}\mathbf{W}_{K}$ via the query and key parameters matrices $\mathbf{W}_{Q},\mathbf{W}_{K}\in\mathbb{R}^{d\times d_{k}}$. We call $d_{k}$ the key dimension. In our setup, the key dimension is not necessarily equal to the embedding dimension divided by the number of attention heads. The per-token attention logit matrix is the product $\mathbf{A}^{\prime}=\mathbf{Q}\mathbf{K}^{T}$. This yields the unnormalized per-token attention weight matrix via the elementwise formula: $a_{i,j}=\exp(a^{\prime}_{i,j}/\sqrt{d_{k}})$. Note that as we define the output in Equation (3) for one input sequence only, we do not have to be explicit about causal attention.
The next token logit displacement matrix $\mathbf{D}=\mathbf{E}\mathbf{W}_{V}\mathbf{W}_{O}\mathbf{U}$ is formed via the value and output parameter matrices $\mathbf{W}_{V}\in\mathbb{R}^{d\times d_{v}}$, $\mathbf{W}_{O}\in\mathbb{R}^{d_{v}\times d}$. We call $d_{v}$ the value dimension. In our setup, the value dimension is not necessarily equal to the key dimension, nor is it necessarily equal to the embedding dimension divided by the number of attention heads.
@@ 4.1 A Hardcoded, Minimal Solution
For any vocabulary size $v$, we now give a formula for a model that is arbitrarily close to being perfect if we choose $C>0$ large enough. It uses embedding and value dimensions $d=d_{v}=v-1$, and key dimension $d_{k}=1$.
We can use the parameter matrices $\mathbf{E}=(I|-\boldsymbol{1})^{T}$, $\mathbf{U}=(-I|\boldsymbol{0})$, $\mathbf{W}_{Q}=\mathbf{W}_{K}=\boldsymbol{0}$, $\mathbf{W}_{V}=vC\mathbf{I}$, and $\mathbf{W}_{O}=I$.
Note that with these parameters, we get
@@
Note that the hardcoded model $f_{\theta}$ has constant attention. Since when training transformers with weight decay induces low-rank attention logit matrices (Kobayashi et al., 2024), we will continue our theoretical investigation with the assumption of constant attention. Therefore, the formula (3) for the next token logit vector $f_{\theta}(\mathbf{t})$ at the input sequence $\mathbf{t}=(t_{1},\dotsc,t_{s})\in\mathbb{X}$ simplifies to the following:
@@ 4.2 Length Generalization at the Price of Less Precision
See Appendix A for the Proof.
4.3 Post-Norm can Resharpen Attention
$f_{\theta}((t_{1}))$ $=\mathbf{B}_{t_{1},:}+\mathbf{D}_{t_{1},:}$
$\text{ and }f_{\theta}((t_{1},t_{2}))$ $=\mathbf{B}_{t_{2},:}+\frac{\mathbf{D}_{t_{1},:}+\mathbf{D}_{t_{2},:}}{2},$
Note that the dilution comes from the multiplier $\frac{1}{s}$ by sequence length. Therefore, we can resharpen attention by applying an RMSNorm (Zhang & Sennrich, 2019) operation to the output. We will verify that this operation indeed fosters length generalization in the experimental, second half of our paper, both in case of the Set Complement Task, and further formal language tasks.
4.4 Training: NLL of One-Hot Sampled Target Distribution
Figure 1: Distribution of the validation TVD at the best checkpoint of each of the 1000 $d=8$, $d_{k}=1$, $d_{v}=8$ minimal transformers trained on the Set Complement Task with ambient set size $v=9$ and varying input sequence length $s$. In each violin plot, the left half shows the values without BEMA, and the right half with BEMA.
4.5 Mitigating Slowdown from Noisy Sampled Targets with BEMA
Figure 2: Distribution of the validation accuracies at the best checkpoints of each of the 100 single attention head architectures trained on the Maximum Retrieval Task with training input sequence lengths sampled uniformly from the set $\{5,\dotsc,16\}$.
We hypothesize that Exponential Moving Average (EMA), a general remedy for gradient noise-induced slowdown, may provide a mitigation in this setting. We use Bias-Corrected Exponential Moving Average (BEMA) (Block & Zhang, 2025), that we now introduce for completeness:
$\theta^{\mathrm{EMA}}_{0}$ $=\theta_{0}\text{ and }$
We call $\alpha_{n}$ a BEMA weight.
5 Experimental Setup
In this Section, we detail the experimental setup in which we verify if the tools we propose indeed foster length generalization.
5.1 Normalization Options
In our survey of related work in Section 2, we in part discuss how normalization layers can be used in transformer models. Following (Kim et al., 2025), in our experiments, we consider the following four normalization options:
pre-norm
Normalize the input of the residual blocks.
post-norm
Normalize residual stream after each residual update.
peri-norm
Normalize the input and the output of each residual block, and the input of the unembedding.
peri-init-norm
Besides the positions listed in peri-norm, also normalize the output of the embedding.
5.2 Hyperparameter Distributions
As we consider various architectural options and various datasets, in order to have a robust comparison, we need to run experiments with various hyperparameter configurations. See for example (Zhang et al., 2022) on the importance of tuning the first and second moment decay rates of Adam (Kingma & Ba, 2017). See Appendix B for the hyperparameter distribution.
5.3 Main Metric: Total Variation Distance
Let $\mathbf{t}\in\mathbb{X}$ be an input sequence. Recall that $f_{\theta}(\mathbf{t})\in\mathbb{R}^{v}$ is the vector of predicted unnormalized next token logits to follow $\mathbf{t}$, and we let $p_{\theta}(\mathbf{t})=\mathop{\mathrm{softmax}}(f_{\theta}(\mathbf{t}))$ denote the corresponding next token probabilities.
The most important metric which we utilize to measure how closely the predicted distribution $p_{\theta}(\mathbf{t})$ approximates the target distribution $p^{*}(\mathbf{t})$, in particular the uniform distribution on legal tokens in case of the Set Complement task (see Equation (1)), is total variation distance (TVD):
$\mathop{\mathrm{TVD}}(\mathbf{t};\theta)=\frac{1}{2}\sum_{t=1}^{v}|p_{\theta}(\mathbf{t})_{i}-p^{*}(\mathbf{t})_{i}|.$ (7)
In further tasks, we use additional metrics, which are discussed as needed.
5.4 Dataloaders, Loss and Metric Aggregation
Both our training and validation dataloaders output minibatches of sequences $\mathbf{T}=(\mathbf{t}_{1},\dotsc,\mathbf{t}_{N})\in\bar{\mathbb{X}}^{N}$. The difference is in the length of the sequences: in order to study length generalization, the training dataloaders output shorter minibatches than the validation dataloaders.
Let $\mathbf{T}$ denote a training or validation minibatch. We follow the standard convention to aggregate the loss and the metrics not only by averaging across the minibatch entries, but also across the sequential dimension. Sometimes, we only aggregate over bins of sequence lengths, or report metrics on given sequence lengths.
5.5 Training
Figure 3: Distribution of the TVD at the best checkpoints of each of the 100 2-layer, 64-dimensional, 4-headed transformers on the $\mathrm{Dyck}(8,10)$ with maximum training input sequence length 700.
Following standard conventions, we initialize parameter matrices with normal distribution of std $\sigma=0.02$ and truncated at $2\sigma$. We use the AdamW (Loshchilov & Hutter, 2019) optimizer. We follow the standard practice of disabling weight decay on embedding, and norm vectors; the latter decision is ablated in (D’Angelo et al., 2024). For learning rate schedule, we use linear warmup, and linear decay (Bergsma et al., 2025). For each of our models, we report two sets of metrics: one for the training parameters $\theta$, and one for the BEMA parameters $\theta^{\mathrm{BEMA}}$, see Subsection 4.5.
6 Experiment Results
6.1 Set Complement Task
Figure 4: Distribution of the ROC AUC scores at tokens 1 resp. 0 at the best checkpoints of each of the 100 2-layer, 64-dimensional, 4-headed transformers on the Tomita 3 resp. 7 languages with maximum training input sequence length 100.
For ambient set sizes $v=9,16$ and input sequence lengths $s=1,\dotsc,8$, we train minimal transformers (see Subsection 3.2) with the minimal possible dimension d=v11d=v-11, $d_{k}=1$, $d_{v}=v-1$ permitted by Theorem 4.2, for 1 000 000 training steps, on minibatches of size $N=256$.
See Figure 1 for the distribution of the best TVD of all our models in case $v=9$. One can see that all normalization methods that include some sort of post-normalization can length generalize from training input sequences of length $s\geq 3$, peri-normalization can robustly length generalize from $s\geq 2$ in concert with our Theorem 4.2, and moreover, it can even length generalize to some extent with training input sequence length 1, thus overcoming our Theorem, that did not suppose post-normalization. We can also see that in most cases, BEMA can boost the best results a little further. See Appendix C.1 for similar results in case $v=16,25$.
6.2 Maximum Retrieval
This task was designed in (Veličković et al., 2025) to showcase the attention dispersion effect in a simple setting. We sample $s$ set entries with a priority $\rho_{i}\sim\mathcal{U}([0,1))$ and a class $\kappa_{i}\sim\mathcal{U}(\{1,\dotsc,10\})$. We get input features $\mathbf{x}_{i}=(\rho_{i}|e_{\kappa_{i}})\in\mathbb{R}^{11}$. These are first embedded via a 2-layer MLP $\psi$ into a hidden representation $\mathbf{h}_{i}\in\mathbb{R}^{128}$. Then attention logits are formed as scalar products [^1] Using a parameter vector $\mathbf{q}$ instead of another 2-layer MLP image $\psi_{q}(q)$ of a random value is our change. $e_{i}=\mathbf{q}\cdot\left(\mathbf{K}\mathbf{h}_{i}\right)^{T}$. These in turn give aggregated values $\mathbf{z}=\sum_{i=1}^{s}\mathrm{softmax}(e_{j})_{i}\mathbf{V}\mathbf{h}_{i}$, which are then transformed to a logit $\mathbf{y}=\phi(\mathbf{z})\in\mathbb{R}^{10}$ by another 2-layer MLP. The model is trained to predict the class $i$ of the highest priority $\rho_{i}$.
We train 100 models on 100 000 minibatches of 128 input sequences of lengths [^2] In the original, the lengths in each minibatch are constant; we removed this restriction. between 5 and 16 and then validated on longer sequences. We test the effect of including a normalization layer after the embedding and after value aggregation, respectively called pre- and post-normalization. We call using normalization at both positions peri-normalization. On Figure 2 we can see that we more robustly get higher accuracies with post- and peri-normalization, and the effect increases with sequence length.
6.3 Bounded Dyck Languages
In the Dyck language $\mathrm{Dyck}(k)$ with $k$ types, we have $k$ pairs of opening $(_{i}$ and closing $)_{i}$ parentheses. Then the valid sequences are the correctly nested sequences. In the theory of formal languages, these languages have central importance by the Chomsky–Schützenberger Representation Theorem (Chomsky & Schützenberger, 1959).
In the bounded Dyck language $\mathrm{Dyck}(k,m)$ of maximum depth $m$, there can be at most $m$ unclosed brackets at any time. It turns out that if we wanted to model hierarchical structures of clauses in “Standard Average European” languages, we could let $m=3$ (Karlsson, 2007).
We include Beginning of Sequence (BOS) and End of Sequence (EOS) tokens in the vocabulary. The first represents an empty sequence, and the latter a complete one. Note that if we train a generative model on random valid sequences, the predicted EOS logit can be used as a complete sequence classification signal.
Note that if we gave each valid token equal probability, then we would usually be in the depths m1m-1 and $m$, making EOS tokens very rare. Therefore, we follow prior work (Hewitt et al., 2020) on benchmarking generative models on learning $\mathrm{Dyck}(k,m)$ in giving the depth increase and depth decrease / EOS events equal probability, thus making opening brackets have probability $\frac{1}{2k}$ when they are legal.
We train 100 2-layer transformers with 4 attention heads, 64 embedding dimensions, and Rotary Positional Encoding (RoPE) (Su et al., 2024) on the task, for 100 000 minibatches of size 64. In Figure 3, we can see that, here too, normalization schemes that include post normalization have better length generalization results, while BEMA can boost the best metrics a little further. See Appendix C.2 for further results.
6.4 Tomita Languages
The Tomita languages are 7 formal languages on the vocabulary $\mathbb{V}=\{\mathrm{[BOS]},0,1,\mathrm{[EOS]}\}$ created to benchmark how well a specific hill-climbing algorithm can find finite state automata that can recognize the languages (Tomita, 1982). As the problem set features various long-dependency relationships, it has become an often used benchmark in training language models to recognize formal languages (Schmidhuber et al., 2001).
Here, we consider languages 3 and 7. In 3, “odd number of consecutive 1s are always followed by an even number of 0s”, while 7 “has the regular expression $0^{*}1^{*}0^{*}1^{*}$(Bhattamishra et al., 2020). Note that if tokens 0, 1, and EOS had the same probability when legal, the sequences would terminate very early. So we set up the probabilities of EOS, and switches between 0 and 1, respectively, so that empirically about half of the sequences terminate by length 100.
Since, potentially for the very skewed token distributions, the TVD results do not give much signal, we show in Figure 4 the ROC AUC values of the tokens 1 and 0, respectively. That is, at each sequence length separately, we collect all validation entry pairs where the token is valid in one and not in the other, and measure the proportion of times that in the valid case the predicted token logit is higher. We can see that the peri-normalization methods are better on validation sequence lengths up to 460 on Tomita 3, while in other cases, they all produce similar results.
7 Conclusion
Length generalization, the ability of performing a task on contexts longer than those seen during training is an essential trait of autonomous agents. Transformers, the leading architecture of our time, usually use softmax attention, which is proven to exhibit attention dispersion, thus performance degrades with sequence length. Via a mechanistic study on the Set Complement Task, we discover that post-normalization, that is normalizing the output of attention blocks may mitigate this effect. We provide experimental evidence on various algorithmic tasks.
Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proof of Theorem 4.2